OpenStudy (anonymous):

find the x and y intercept -4x^2+15x+25=0

OpenStudy (amistre64):

the y intercept is when x = 0; which makes this (0,25)

OpenStudy (amistre64):

the x intercept is when this equation = 0

OpenStudy (amistre64):

-(4x^2 -15x -25) = 0 we need 2 numbers (b and c) that subtract to get 15 and multiply to get -100

OpenStudy (anonymous):

http://myalgebra.com/algebra_solver.aspx go here and you'll get it

OpenStudy (amistre64):

100 factors into: 1,100 2,50 4,25 5,20 .... 5-20 = -15; and 5(-20) = -100 so those fit the bill

OpenStudy (amistre64):

(x-20/4) (x+5/4) when x = 5 and x = -5/4 look to be the x intercepts